# Contract Design for Pandora’s Box

### v1

@@ 4.2 Binary Boxes
We study a subclass of the problem where every box $i\in[n]$ contains two prizes with positive probability. One of those prizes has value 0 for both $\mathcal{P}$ and $\mathcal{A}$, and is called 0-prize. The other prize (the positive prize) is denoted by the value pair $(a_{i},b_{i})$ with $a_{i},b_{i}\geq 0$ and has probability $p_{i}\in(0,1]$. Consequently, the 0-prize is drawn from box $i$ with probability $1-p_{i}$. Note that there cannot be a positive payment for the 0-prize. Thus, payments for box $i$ are given by a single $t_{i}\in[0,b_{i}]$, which specifies the payment for the positive prize.
@@
For the two remaining cases ($i^{*}>i$ and $i^{*}<i$) we show that the optimal ordering can be changed sequentially into one that has $i^{*}=i$ without decreasing the expected value for $\mathcal{P}$. We discuss the proof for case 1: $i^{*}>i$. The proof of the other case is very similar and deferred to Appendix B.
@@
$p_{k}(b_{k}-\Delta_{i_{r}}-\Delta_{i_{n},i_{r}})+(1-p_{k})\smash{\tilde{p_{1}}}\smash{\tilde{b_{1}}}\;\geq\;\smash{\tilde{p_{1}}}\smash{\tilde{b_{1}}}+(1-\smash{\tilde{p_{1}}})p_{k}(b_{k}-\Delta_{i_{r}})$
@@
$b_{k}\geq\smash{\tilde{b_{1}}}+\Delta_{i_{r}}+\Delta_{i_{n},i_{r}}/\smash{\tilde{p_{1}}}$
@@
>\;> $\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+(1-\smash{\tilde{p_{2}}})p_{k}(b_{k}-\Delta_{i^{*}_{n}})+(1-\smash{\tilde{p_{2}}})(1-p_{k})p_{r}(b_{r}^{k}-\Delta_{i^{*}_{n}})$
@@
and since $b_{r}^{k}<b_{k}$
$b_{r}^{k}>\smash{\tilde{b_{2}}}+\Delta_{i^{*}_{n}}+\Delta_{i^{*}_{n},i_{r}}/\smash{\tilde{p_{2}}}$ (3)
@@
>\;> $p_{k}(b_{k}-\Delta_{i^{*}_{n}}-\Delta_{i^{*}_{n},i_{r}})+(1-p_{k})p_{r}(b_{r}^{k}-\Delta_{i^{*}_{n}}-\Delta_{i^{*}_{n},i_{r}})+(1-p_{k})(1-p_{r})\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}$
@@
$(-p_{r}-\smash{\tilde{p_{2}}}+p_{r}\smash{\tilde{p_{2}}})b_{k}+(1-p_{r})\smash{\tilde{p_{2}}}\Delta_{i^{*}_{n}}\;>\;-(1-p_{r})\Delta_{i_{n}^{*},i_{r}}-p_{r}b_{r}^{k}-\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+p_{r}\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}$
@@
$(1-p_{r})\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+(1-p_{r})\smash{\tilde{p_{2}}}\Delta_{i^{*}_{n}}+(1-p_{r})\Delta_{i_{n}^{*},i_{r}}\;>\;\smash{\tilde{p_{2}}}(1-p_{r})b_{k}+p_{r}(b_{k}-b_{r}^{k})$
This implies $p_{r}<1$, since otherwise we derive $b_{k}-b_{r}^{k}<0$, a contradiction. Now if $p_{r}\in(0,1)$, $\smash{\tilde{p_{2}}}\in(0,1]$, and $b_{r}^{k}<b_{k}$, the above implies
$\smash{\tilde{b_{2}}}+\Delta_{i^{*}_{n}}+\Delta_{2}/\smash{\tilde{p_{2}}}\;$ $>\;b_{k}$ (4)
The equations (3) and (4) imply a contradiction with $b_{r}^{k}<b_{k}$.
@@
4.3 I.I.D. with Single Positive Prize for Principal
Theorem 4.
Lemma 6.
Example 1.
Now suppose for some $k\in\{1,\ldots,n-1\}$
$\alpha=\frac{(1-1/n)^{k}}{1-(1-1/n)^{k}}\enspace.$
Lemma 7.
During Phase 2, the process only stops early when an an outcome $j>j^{*}$ is drawn. In this case, the utility for $\mathcal{P}$ is 0. If the process does not stop early, $\mathcal{P}$ receives a utility only if outcome 0 is the best outcome for the agent among all $n$ outcomes.
  • in all remaining boxes, some outcome $1\leq j^{\prime}<j$ was drawn.
Thus, the expected utility for $\mathcal{P}$ for Phase 2 is
With this at hand, we show the following lemma.
Lemma 8.
Proof.
Consider an optimal contract with multiple outcomes that are targeted with the payment for outcome 0. Then there are $j_{1},j_{2}\in[j^{*}]$ with $j_{1}<j_{2}$ such that $n_{j_{1}}>0$ and $n_{j_{2}}>0$. Assume w.l.o.g. that the outcomes between $j_{1}$ and $j_{2}$ are never targeted, i.e., $n_{j_{1}+1}=\dots=n_{j_{2}-1}=0$. Suppose we could swap a $\delta$-fraction of boxes from target outcome $j_{2}$ to $j_{1}$ (or vice versa), for some $\delta\in\mathbb{R}$. Then let $n_{j_{1}}^{\prime}:=n_{j_{1}}+\delta$, $n_{j_{2}}^{\prime}:=n_{j_{2}}-\delta$, $n_{j}^{\prime}=n_{j}$ for all $j^{\prime}\in[j^{*}]\setminus\{j_{1},j_{2}\}$. Accordingly, define $N_{j}^{\prime}:=\sum_{j^{\prime}=1}^{j}n_{j^{\prime}}^{\prime}$. We want to maximize the expected utility for $\mathcal{P}$ over $\delta$. It holds that
[…]
We used that for all $j\in\{j_{1}+1,\dots,j_{2}-1\}$, it holds that $n_{j}^{\prime}=0$ and therefore $N_{j}^{\prime}=N_{j-1}^{\prime}$. For any extreme point, it follows that
Considering the second derivative at any extreme point, it holds that
[…]
The last inequality holds because $j_{1}<j_{2}$, so $Q_{j_{1}}<Q_{j_{2}}$, and $\frac{p}{Q_{j_{2}}}>\frac{p}{Q_{j_{2}}}$. Furthermore, $\ln\left(\frac{p}{Q_{j_{1}}}+1\right)>\ln\left(\frac{p}{Q_{j_{2}}}+1\right)\geq 0$ due to monotonicity of the logarithm and $\frac{p}{Q_{j_{2}}}\geq 0$. All other terms of the product are also non-negative. As the second derivative is positive at every extremum, there can be at most one local minimum, but no local maximum. Hence, we maximize the function by choosing either the smallest or largest possible value for $\delta$. Therefore, it represents an improvement to either decrease $n_{j_{1}}$ to 0 and increase $n_{j_{2}}$ by $n_{j_{1}}$, or vice versa. Applying this insight repeatedly, the statement of the lemma follows. ∎
@@ Appendix A Missing Material from Section 2
Thus, in the remainder of this proof, we consider the boxes with strictly positive cost. We partition these boxes into phases according to their fair caps. Consider a phase in which all fair caps equal $\smash{\varphi}$. At the beginning of each phase, when $\mathcal{A}$ opens the first box, there might be an outcome $j\in[m]$ from a box $i$ from a former phase with $a_{ij}+t_{ij}\leq\smash{\varphi}$ with largest utility (for $\mathcal{A}$) but not large enough to stop in a former phase. During a phase with fair cap $\smash{\varphi}$, $\mathcal{A}$ clearly has to stop if an outcome $j$ with $a_{ij}+t_{ij}>\smash{\varphi}$ has been drawn from a box $i$, in order to be consistent with the index policy. Also, whenever an outcome $j$ with $a_{ij}+t_{ij}<\smash{\varphi}$ is drawn from a box $i$, (and the next box belongs to the current phase, i.e., has fair cap $\smash{\varphi}$), then $\mathcal{A}$ has to continue. From $\mathcal{P}$’s point of view, for each box we can summarize the outcomes where $\mathcal{A}$ needs to stop in a single outcome (with combined probability and utility equal to conditional expectation of the utility). For the sake of the current phase, all outcomes where $\mathcal{A}$ needs to continue can also be summarized in the same way.
Thus, for each phase, we can formulate the problem of finding a policy that is optimal for $\mathcal{P}$ (subject to being optimal for $\mathcal{A}$) as follows: There are $n$ boxes and each one contains one of $m$ possible outcomes for $\mathcal{P}$, which is drawn independently from a known distribution. For each box $i\in[n]$, let $m_{i}$ denote the number of outcomes with positive probability, and the non-negative values for $\mathcal{P}$ are given by $b_{i,1},\dots,b_{i,m_{i}}$ (where we re-use $b$ for the ease of notation), with respective probabilities $p_{i,1},\dots,p_{i,m_{i}}$. We assume w.l.o.g. that $b_{i,2}\geq b_{i,3}\geq\dots\geq b_{i,m_{i}-1}\geq b_{i,m_{i}}=0$. Opening boxes has no cost, but whenever outcome 1 is drawn from some box $i$, the process stops immediately with value $b_{i,1}$. Moreover, box $i$ has to be opened when in every previous round $i^{\prime}$, outcome $m_{i^{\prime}}$ was drawn. In all other cases, i.e., if $j(i^{\prime})\in\{2,\dots,m_{i}^{\prime}-1\}$ was the outcome drawn from box $i^{\prime}$, for all $i^{\prime}<i$, then $\mathcal{P}$ can decide whether to stop, yielding $\max_{i^{\prime}<i}b_{i^{\prime},j(i^{\prime})}$, or continue with opening box $i$.
For every box $i\in[n]$, let
$\tau_{i}:=\max_{k\in[m_{i}-1]}\frac{\sum_{i=1}^{k}p_{i}b_{i}}{\sum_{i=1}^{k}p_{i}},$
@@
i.e., $\tau_{i}$ is the maximum expected value for $\mathcal{P}$ for an outcome from box $i$ conditioned on the event that an outcome from that box is taken (subject to being consistent with the index policy). Consequently, $\tau_{i}\geq b_{i,1}$.
@@
An equivalent formulation of the problem is the following: (Weakly) increase the value of outcome $1$ to $b_{i,1}^{\prime}:=\tau_{i}$, but, whenever outcome 1 is drawn, costs of $c_{i,1}:=\tau_{i}-b_{i,1}\geq 0$ are incurred. All other outcomes keep their value and no costs occur for them, i.e., for $j>1$, set $b_{i,j}^{\prime}:=b_{i,j}$ and $c_{i,j}:=0$. Similarly as in [31], to analyze the (expected) performance of any policy for this problem, let
@@
$A_{i}$ $:=\mathbf{1}_{\text{Outcome }J_{i}\text{ from box }i\in[n]\text{ is taken}}.$
We can only select outcome $J_{i}$ from box $i$ if the box was opened before:
@@
Furthermore, if box $i\in[n]$ is opened and contains outcome 1, then we have to stop and select it:
$I_{i}=1\text{ and }J_{i}=1\implies A_{i}=1\quad\text{for all}~{}i\in[n].$ (6)
@@
$u_{\mathcal{P}}=\sum_{i=1}^{n}A_{i}\cdot b_{i,J(i)}=\sum_{i=1}^{n}A_{i}\cdot\left(b^{\prime}_{i,J(i)}-c_{i,J(i)}\right).$
@@
$\mathbb{E}[u_{\mathcal{P}}]$ $=\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}p_{i,j}\cdot A_{i}\cdot\left(b^{\prime}_{i,j}-c_{i,j}\right)$
$=\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}p_{i,j}\cdot A_{i}\cdot b^{\prime}_{i,j}-\sum_{i=1}^{n}p_{i,1}\cdot A_{i}\cdot c_{i,1}$
$=\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}p_{i,j}\cdot A_{i}\cdot b^{\prime}_{i,j}-\sum_{i=1}^{n}p_{i,1}\cdot I_{i}\cdot c_{i,1}.$ (7)
The last step holds because, whenever $J(i)=1$, (5) and (6) imply $A_{i}=I_{i}$.
@@
We now create an instance for the (original) Pandora’s Box problem. We keep values $b^{\prime}_{i,1}\dots,b^{\prime}_{i,m_{i}}$ and probabilities $p_{i,1},\dots,p_{i,m_{i}}$ for every box $i\in[n]$. Moreover, we have the opening costs of $c_{i}:=p_{i,1}\cdot c_{i,1}$ for every box $i\in[n]$. Recall that, in Pandora’s Box problem there are no costs for certain outcomes or obligations to accept them. In particular, if we define $\tilde{A}_{i}$ and $\tilde{I}_{i}$ as the analogs of the above indicator variables for a policy for the Pandora’s Box problem, then the analog of (6) is not required. The utility that the policy obtains in expectation is on the other hand
$\sum_{i=1}^{n}\mathbb{E}[\tilde{A}_{i}\cdot b^{\prime}_{i,J(i)}-\tilde{I}_{i}\cdot c_{i}]=\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}p_{i,j}\cdot\tilde{A}_{i}\cdot b^{\prime}_{i,j}-\sum_{i=1}^{n}\tilde{I}_{i}\cdot c_{i},$ (8)
@@
cf. [31], which is the analog of (7). As such, the Pandora’s Box instance is a relaxation of the former instance. There does, however, exist an optimal policy for the Pandora’s Box instance such that $\tilde{A}_{i}$ and $\tilde{I}_{i}$ fulfill the analog of (6): The fair cap of $\smash{\varphi_{i}}$ is $\tau_{i}$ for every $i\in[n]$. Thus $b^{\prime}_{1,i}=\smash{\varphi_{i}}$ for all $i\in[n]$, and the index policy always stops and selects outcome $1$ when it is realized.
@@
Therefore, the following policy is optimal for the former problem (with the obligation to stop whenever outcome 1 is observed): Consider boxes in the order of $\tau_{i}$. When box $i$ is considered and outcome 1 was not drawn before, open box $i$ unless $\max_{i^{\prime}<i}b_{i^{\prime},j(i^{\prime})}\geq\tau_{i}$. Always stop when outcome 1 is drawn. Break ties arbitrarily. ∎
@@
Appendix C Missing Material from Section 4.3
Proof of Lemma 6..
@@
Proof of Lemma 7..

### v2

@@ 4.2 Binary Boxes
We study a subclass of the problem where every box $i\in[n]$ contains two prizes with positive probability. One of those prizes has value 0 for both $\mathcal{P}$ and $\mathcal{A}$, and is called 0-prize. The other prize (the positive prize) is denoted by the value pair $(a_{i},b_{i})$ with $a_{i},b_{i}\geq 0$ and has probability $p_{i}\in(0,1]$. Consequently, the 0-prize is drawn from box $i$ with probability $1-p_{i}$. Note that there cannot be a positive payment for the 0-prize. Thus, payments for box $i$ are given by a single $t_{i}\geq 0$, the payment for the positive prize.
@@
For the two remaining cases ($i^{*}>i$ and $i^{*}<i$) we show that the optimal ordering can be changed sequentially into one that has $i^{*}=i$ without decreasing the expected value for $\mathcal{P}$. We discuss the proof for case 1: $i^{*}>i$. The proof of the other case is very similar (see Appendix B).
@@
$p_{k}(b_{k}-\Delta_{i_{r}}-\Delta_{i_{n},i_{r}})+(1-p_{k})\smash{\tilde{p_{1}}}\smash{\tilde{b_{1}}}\;\geq\;\smash{\tilde{p_{1}}}\smash{\tilde{b_{1}}}+(1-\smash{\tilde{p_{1}}})p_{k}(b_{k}-\Delta_{i_{r}}),$
@@
$b_{k}\geq\smash{\tilde{b_{1}}}+\Delta_{i_{r}}+\Delta_{i_{n},i_{r}}/\smash{\tilde{p_{1}}}.$
@@
>\;> $\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+(1-\smash{\tilde{p_{2}}})p_{k}(b_{k}-\Delta_{i^{*}_{n}})+(1-\smash{\tilde{p_{2}}})(1-p_{k})p_{r}(b_{r}^{k}-\Delta_{i^{*}_{n}}).$
@@
and, since $b_{r}^{k}<b_{k}$,
$b_{r}^{k}>\smash{\tilde{b_{2}}}+\Delta_{i^{*}_{n}}+\Delta_{i^{*}_{n},i_{r}}/\smash{\tilde{p_{2}}}.$ (3)
@@
>\;> $p_{k}(b_{k}-\Delta_{i^{*}_{n}}-\Delta_{i^{*}_{n},i_{r}})+(1-p_{k})p_{r}(b_{r}^{k}-\Delta_{i^{*}_{n}}-\Delta_{i^{*}_{n},i_{r}})+(1-p_{k})(1-p_{r})\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}.$
@@
$(-p_{r}-\smash{\tilde{p_{2}}}+p_{r}\smash{\tilde{p_{2}}})b_{k}+(1-p_{r})\smash{\tilde{p_{2}}}\Delta_{i^{*}_{n}}\;>\;-(1-p_{r})\Delta_{i_{n}^{*},i_{r}}-p_{r}b_{r}^{k}-\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+p_{r}\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}},$
@@
$(1-p_{r})\smash{\tilde{p_{2}}}\smash{\tilde{b_{2}}}+(1-p_{r})\smash{\tilde{p_{2}}}\Delta_{i^{*}_{n}}+(1-p_{r})\Delta_{i_{n}^{*},i_{r}}\;>\;\smash{\tilde{p_{2}}}(1-p_{r})b_{k}+p_{r}(b_{k}-b_{r}^{k}).$
This implies $p_{r}<1$, since otherwise we derive $b_{k}-b_{r}^{k}<0$, a contradiction. Now, if $p_{r}\in(0,1)$, $\smash{\tilde{p_{2}}}\in(0,1]$, and $b_{r}^{k}<b_{k}$, the above implies
$\smash{\tilde{b_{2}}}+\Delta_{i^{*}_{n}}+\Delta_{2}/\smash{\tilde{p_{2}}}\;$ $>\;b_{k}.$ (4)
Equations (3) and (4) imply a contradiction with $b_{r}^{k}<b_{k}$.
@@ Appendix A Missing Material from Section 2
Thus, in the remainder of this proof, we consider the boxes with strictly positive cost. We partition these boxes into phases according to their fair caps, i.e., all boxes with an identical fair cap belong to the same phase. The optimal order of phases is unambiguously determined by the index policy. Hence with respect to optimal tie-breaking, considering the $\mathcal{P}$-optimal action within each phase separately, conditioned on the previous realizations, suffices.
Let $\smash{\varphi}$ be the fair cap of any phase. If we consider the first phase, we define $a^{*}=0$. Otherwise, let option $j^{\prime}$ from box $i^{\prime}$ be a best option (for $\mathcal{A}$) from previous phases, with ties broken in favor of $\mathcal{P}$. Then $a^{*}:=a_{i^{\prime}j^{\prime}}+t_{i^{\prime}j^{\prime}}$ is the utility for $\mathcal{A}$ for this option. Note that the current phase is only reached when $a^{*}\leq\smash{\varphi}$, and does not need to be considered otherwise.
For each box $i$ of the current phase, we partition the set of options with positive probability depending on the behavior of $\mathcal{A}$ during the phase, which is given by $\mathcal{A}$’s index policy. Let
$\Omega_{i}^{>}:=\{j\in[m]~{}:~{}p_{ij}>0,~{}a_{ij}+t_{ij}>\smash{\varphi_{i}}\}$
denote the set of options from box $i$ where $\mathcal{A}$ would stop immediately. Furthermore, let
$\Omega_{i}^{<}:=\{j\in[m]~{}:~{}~{}p_{ij}>0,~{}a_{ij}+t_{ij}<\smash{\varphi_{i}}\}$
denote the set of options from box $i$ that $\mathcal{A}$ would never take immediately (unless all boxes of the current phase are opened). Upon observing all other options, $\mathcal{A}$ is indifferent, and they are collected in the set
$\Omega_{i}^{=}:=\{j\in[m]~{}:~{}p_{ij}>0,~{}a_{ij}+t_{ij}=\smash{\varphi_{i}}\}.$
Note that if $a^{*}=\smash{\varphi}$, $\mathcal{A}$ can stop at any time and take the previously revealed value of $a^{*}$. Additionally, note that the set $\Omega_{i}^{>}$ is never empty, as boxes have strictly positive opening costs.
We reformulate the problem of finding a $\mathcal{P}$-optimal policy among all $\mathcal{A}$-optimal policies for the current phase as follows: There are $n$ boxes in the current phase, and there is a fallback option with value $v^{*}$, where
$v^{*}=\begin{cases}0,&\text{ if }a^{*}<\smash{\varphi},\\ b_{i^{\prime}j^{\prime}}-t_{i^{\prime}j^{\prime}},&\text{ otherwise.}\end{cases}$
For each box $i\in[n]$, there are $m_{i}:=|\Omega_{i}^{=}|+2$ options that are drawn independently from a known distribution. Option 1 has probability $\tilde{p}_{i,1}$ and value $v_{i,1}$, and combines all options from $\Omega_{i}^{>}$ that make $\mathcal{A}$ stop immediately. The value is the conditioned expected utility of those options for $\mathcal{P}$. More formally,
$\tilde{p}_{i,1}:=\sum_{j\in\Omega_{i}^{>}}p_{ij}\quad\text{and}\quad v_{i,1}:=\frac{\sum_{j\in\Omega_{i}^{>}}(b_{ij}-t_{ij})\cdot p_{ij}}{\sum_{j\in\Omega_{i}^{>}}p_{ij}}.$
Furthermore, the options $12,\dots,m_{i}-$ 1 are given by the set $\Omega_{i}^{=}$. Their probabilities $\tilde{p}_{i,j}$ are identical to to their respective counterpart in $\Omega_{i}^{=}$, while their values $v_{i,j}$ are given by the respective utility for $\mathcal{P}$. We assume w.l.o.g. that $v_{i,2}\geq v_{i,3}\geq\dots\geq v_{i,m_{i}-1}\geq 0$. Finally, the last option $m_{i}$ combines all options in $\Omega_{i}^{<}$ that $\mathcal{A}$ will never take before having inspected all $n$ boxes of the phase. Let
$\tilde{p}_{i,m_{i}}:=\sum_{j\in\Omega_{i}^{<}}p_{ij}\quad\text{and}\quad v_{i,m_{i}}:=0.$
The boxes can be opened in any order without opening costs. Whenever outcome 1 is drawn from some box $i$, the process stops immediately with value $v_{i,1}$. Otherwise, the process can be stopped voluntarily in case there is a strictly positive value available in an opened box or as fallback value $v^{*}$, and it stops after $n$ rounds at the latest. When stopping without having option 1 drawn, we receive the maximum value from all opened boxes (including the fallback value $v^{*}$).
We need to argue why our formulated problem is equivalent to the problem of finding a $\mathcal{P}$-optimal tie-breaking rule. When all boxes of the current phase are opened, the expected utilities are not always preserved: In the original process, options from the set $\Omega_{i}^{<}$ of some box $i$ can possibly yield a positive value for $\mathcal{P}$ after the phase is completed, but it holds $v_{i,m_{i}}=0$ in the reformulated process. Crucially, the values within $\Omega_{i}^{<}$ never affect tie-breaking: In order to obtain a positive value with an option from $\Omega_{i}^{<}$, the following must hold:
  1. 1.
    Every box from the current phase has to be opened.
  2. 2.
    Every box $i^{\prime}$ from the current phase has to contain an option from $\Omega_{i^{\prime}}^{<}$, as any option from $\Omega_{i^{\prime}}^{=}$ or $\Omega_{i^{\prime}}^{>}$ would be preferred by $\mathcal{A}$.
  3. 3.
    $v^{*}=0$, because a fallback option with positive value for $\mathcal{P}$ would be preferred by $\mathcal{A}$.
As all boxes have to be opened, the opening order is irrelevant. All decisions regarding the tie-breaking are made without incorporating the actual value of drawn options from $\Omega_{i^{\prime}}^{<}$ of any box $i^{\prime}$: Whenever $\mathcal{A}$ is indifferent, it must be upon observing an option from $\Omega_{i^{\prime}}^{=}$ after opening some box $i^{\prime}$, or due to a positive fallback value. But then, an option from $\Omega_{i}^{<}$ will never be considered by $\mathcal{A}$.
Therefore, the choices for $\mathcal{P}$-optimal tie-breaking are preserved with our formulated problem. The same is true for the original expected utilities whenever they can influence the optimal action (opening order and stopping behavior). Therefore an optimal action for the formulated problem constitutes an optimal tie-breaking rule.
We apply further adjustments to derive an optimal policy for our formulated problem. For every box $i\in[n]$, let
$\tau_{i}:=\max_{k\in[m_{i}-1]}\frac{\sum_{j=1}^{k}\tilde{p}_{ij}v_{ij}}{\sum_{j=1}^{k}\tilde{p}_{ij}},$
@@
i.e., $\tau_{i}$ is the maximum expected value for $\mathcal{P}$ for an option from box $i$ conditioned on the event that an option from that box is taken (subject to being consistent with $\mathcal{A}$’s index policy). Consequently, $\tau_{i}\geq v_{i,1}$.
@@
An equivalent formulation of the problem is the following: (Weakly) increase the value of option $1$ to $v_{i,1}^{\prime}:=\tau_{i}$, but, whenever option 1 is drawn, costs of $c_{i,1}:=\tau_{i}-v_{i,1}\geq 0$ are incurred. All other options keep their value and no costs occur for them, i.e., for $j>1$, set $v_{i,j}^{\prime}:=v_{i,j}$ and $c_{i,j}:=0$.
Similarly as in [33], to analyze the (expected) performance of any policy for this problem, let
@@
$A_{i}$ $:=\mathbf{1}_{\text{Option }J(i)\text{ from box }i\in[n]\text{ istaken}}.$
We can only select option $J(i)$ from box $i$ if the box was opened before:
@@
Furthermore, if box $i\in[n]$ is opened and contains option 1, then we have to stop and select it:
$I_{i}=1\text{ and }J(i)=1\implies A_{i}=1\quad\text{for all}~{}i\in[n].$ (6)
@@
$u_{\mathcal{P}}=\sum_{i=1}^{n}A_{i}\cdot v_{i,J(i)}=\sum_{i=1}^{n}A_{i}\cdot\left(v^{\prime}_{i,J(i)}-c_{i,J(i)}\right).$
@@
$\mathbb{E}[u_{\mathcal{P}}]$ $=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot\left(v^{\prime}_{i,J(i)}-c_{i,J(i)}\right)\right]$
$=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot v^{\prime}_{i,J(i)}\right]-\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot c_{i,J(i)}\right]$
$=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot v^{\prime}_{i,J(i)}\right]-\sum_{i=1}^{n}\Pr[J(i)=1\cap A_{i}=1]\cdot c_{i,1}$
$=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot v^{\prime}_{i,J(i)}\right]-\sum_{i=1}^{n}\Pr[J(i)=1\cap I_{i}=1]\cdot c_{i,1}$
$=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot v^{\prime}_{i,J(i)}\right]-\sum_{i=1}^{n}\tilde{p}_{i,1}\cdot\Pr[I_{i}=1]\cdot c_{i,1}$
$=\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\cdot v^{\prime}_{i,J(i)}\right]-\mathbb{E}\left[\sum_{i=1}^{n}I_{i}\cdot\tilde{p}_{i,1}\cdot c_{i,1}\right].$ (7)
We use that $c_{i,J(i)}=0$ unless $J(i)=1$, but in this case, (5) and (6) imply $A_{i}=I_{i}$. Then we use that $I_{i}=1$ and $J(i)=1$ are independent events.
@@
We now create an instance for the (original) Pandora’s Box problem. We keep values $v^{\prime}_{i,1}\dots,v^{\prime}_{i,m_{i}}$ and probabilities $\tilde{p}_{i,1},\dots,\tilde{p}_{i,m_{i}}$ for every box $i\in[n]$. Moreover, we have the opening costs of $c_{i}:=\tilde{p}_{i,1}\cdot c_{i,1}$ for every box $i\in[n]$. Recall that, in Pandora’s Box problem there are no costs for certain options or obligations to accept them. In particular, if we define $\tilde{A}_{i}$ and $\tilde{I}_{i}$ as the analogs of the above indicator variables for a policy for the Pandora’s Box problem, then the analog of (6) is not required. The utility that the policy obtains in expectation is on the other hand (c.f. [33])
$\mathbb{E}\left[\sum_{i=1}^{n}\tilde{A}_{i}\cdot v^{\prime}_{i,J(i)}-\tilde{I}_{i}\cdot c_{i}\right]=\mathbb{E}\left[\sum_{i=1}^{n}\tilde{A}_{i}\cdot v^{\prime}_{i,J(i)}\right]-\mathbb{E}\left[\sum_{i=1}^{n}\tilde{I}_{i}\cdot\tilde{p}_{i,1}\cdot c_{i,1}\right],$ (8)
@@
which is the analog of (7). As such, the Pandora’s Box instance is a relaxation of the former instance. There does, however, exist an optimal policy for the Pandora’s Box instance such that $\tilde{A}_{i}$ and $\tilde{I}_{i}$ fulfill the analog of (6): The fair cap $\smash{\varphi_{i}}$ of box $i$ is $\tau_{i}$ for every $i\in[n]$.
This condition can be seen as follows. Recall that by definition, $\tau_{i}=\left(\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v_{ij}\right)/\left({\sum_{j=1}^{k^{*}}\tilde{p}_{ij}}\right)$ for some $k^{*}\in[m_{i}-1]$, such that $v^{\prime}_{i,j}\geq\tau_{i}$ for all $j\in[k^{*}]$, and $v^{\prime}_{i,j}\leq\tau_{i}$ for all $j\in[m_{i}]\setminus[k^{*}]$. Otherwise, $\tau_{i}$ was not chosen as the correct maximum. If $k^{*}=1$, it follows $\tau_{i}=v_{i1}$ and therefore $c_{i}=0$ in our constructed instance. In this case, the fair cap is ambiguous and must only fulfill $\smash{\varphi_{i}}\geq\max_{j\in[m_{i}]}v_{ij}=v_{i1}$, which is true. Otherwise, it holds $\tau_{i}>v_{i1}$ and therefore $c_{i}>0$, and the fair cap is unique. We now prove that $\smash{\varphi_{i}}=\tau_{i}$ fulfills the definition of the fair cap:
$\sum_{j=1}^{m_{i}}\tilde{p}_{ij}\cdot\max\{0,v^{\prime}_{ij}-\tau_{i}\}=c_{i}~{}\Leftrightarrow~{}\sum_{j=1}^{k^{*}}\tilde{p}_{ij}\cdot(v^{\prime}_{ij}-\tau_{i})=c_{i}~{}\Leftrightarrow~{}\frac{\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v^{\prime}_{ij}-c_{i}}{\sum_{j=1}^{k^{*}}\tilde{p}_{ij}}=\tau_{i}.$
This equals the definition of $\tau_{i}$, because in the numerator, we have
$\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v^{\prime}_{ij}-c_{i}=\sum_{j=2}^{k^{*}}\tilde{p}_{ij}v^{\prime}_{ij}+\tilde{p}_{i1}v^{\prime}_{i1}-c_{i}=\sum_{j=2}^{k^{*}}\tilde{p}_{ij}v_{ij}+\tilde{p}_{i1}\tau_{i}-\tilde{p}_{i1}\cdot(\tau_{i}-v_{i1})=\sum_{j=1}^{k^{*}}\tilde{p}_{ij}v_{ij}.$
@@
Thus $v^{\prime}_{i,1}=\smash{\varphi_{i}}$ for all $i\in[n]$, and stopping with option $1$ when it is realized is always in accordance with the index policy. Therefore, the following policy is optimal for the former problem (with the obligation to stop whenever option 1 is observed): Consider boxes in the order of $\tau_{i}$. When box $i$ is considered and option 1 was not drawn before, open box $i$ unless the maximum value from an opened box or the fallback value is greater than $\tau_{i}$. Always stop when option 1 is drawn. Break any remaining ties arbitrarily. ∎
@@
Appendix C I.I.D. with Single Positive Prize for Principal
Theorem 5.
Lemma 5.
Proof.
@@
Example 1.
Now suppose for some $k\in\{1,\ldots,n-1\}$
$\alpha=\frac{(1-1/n)^{k}}{1-(1-1/n)^{k}}\enspace.$
Lemma 6.
Proof.